001 10.0 points A quantum of electromagnetic radiation has an energy of 1.68 keV. What is its wavelength? The speed of light is 2.99792 x 10* m/s, and Plancks constant is 6.62607 x 10 4 J.s Answer in units of nm.
Answer #1

Here,

energy of a quantum = 1.68 keV

for the wavelength

energy of a photon = h* c/wavelength

1.68 *10^3 * 1.602 *10^-19 = 6.626 *10^-34 * 2.9972 *10^8/wavelength

wavelength = 7.3789634* 10^-10 m

the wavelength is 7.3789634* 10^-10 m